e The Future Worth in Year 10 x xle For The Cash Flow x G ion /connect.html er 2 homework Problem 02.030- Arithmetic change in cash flows sh flow gradient series equals $350 in year 1. $450 in year 2, and amounts increasing by $100 per year through year An arithmetic ca 11. At 1-11% per year, determine the present worth of the cash flow series in year . The present worth of the cash flow series in year O is $ Prey 401 9Hİ Next >
Answer #1

Present worth = (P/A, i , N) + (P/G, i, N)

Present worth= A (P/A, 11%,11) + G (P/G, 11%,11)

Present worth =350(P/A, 11%,11) + 100(P/G, 11%,11)

Present worth =350(6.2065153253524) +100(24.69453517299)

Present worth =2172.28 + 2469.454= 4641.734

The present worth of the cash flow series in year 0 is $4641.734


7 J, YI